The length of a rocket is measured to be 6.0 m by an observer on Earth as it moves with respect to the Earth at a speed of 0.8c. What is the length of the rocket as measured by an astronaut inside the rocket?
1
Expert's answer
2019-01-28T15:58:43-0500

For the "outside" observer the length

"L"

of a moving rocket is

"L=L_0\cdot \sqrt{1-v^2/c^2},"

where

"L_0"

- the length measured in the inertial frame by an observer resting relative to the rocket. Thus

"L_0=\frac{L}{\sqrt{1-v^2/c^2}}=\frac{6}{\sqrt{1-(0.8c)^2/c^2}}=10 \text{ m}."